# Koopman for stochastic dynamics: error bounds for kernel extended dynamic mode decomposition

### v1

@@ 5. Numerical examples
$X_{0}$ $=x_{0}\in\Omega$
with Lipschitz continuous drift $b:\mathbb{R}\to\mathbb{R}$ and diffusion $\sigma:\mathbb{R}\to\mathbb{R}$, and Brownian motion $(W_{t})_{t\geq 0}$, via the Euler-Maruyama method (see, e.g., [KloePlat1992]). [Okse2000, Theorem 5.2.1] guarantees the existence of a unique solution $(X_{t})_{t\geq 0}$. A basic example for this case is the Ornstein-Uhlenbeck process (OU process) given by the linear SDE
@@
We fix a timestep $T>0$, set $X=X_{0}$ and $Y=X_{T}$. To apply the setting of Sections 3 and 4, the solution process should remain in some bounded domain $\Omega=(a,b)$ until time $T$. [PhilSchaWort2025] elaborates on a sufficient condition on the drift and diffusion for the invariance of a compact interval under the SDE flow. Another way to guarantee the boundedness is by reflection at the boundaries. We consider this doubly reflected OU process on $\Omega=(0,1)$. It is formally given by the SDE
@@
where the $L_{t}\geq 0$ and $U_{t}\geq 0$ are the minimal nondecreasing processes that ensure $X_{t}\geq 0$ and $X_{t}\leq 1$ with reflection at the boundaries for $t\geq 0$, respectively. The processes $L$ and $U$ can be interpreted as “idleness process” and “loss process”; they satisfy $\int_{0}^{T}\chi_{\{X_{t}>0\}}\,dL_{t}=0$ and $\int_{0}^{T}\chi_{\{X_{t}<1\}}\,dU_{t}=0$ for any $T>0$, respectively. We refer to [HuanMandSpre2014] for details.
@@
$\mathbf{K}_{\mathcal{X}}=\bigl(k(x_{i},x_{j})\bigr)_{i,j=1}^{d}.$
@@
$(\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}})_{ij}=\frac{1}{m_{\mathrm{train}}}\sum_{m=1}^{m_{\mathrm{train}}}k(x_{i},y_{j}^{(m)}).$
@@
$k_{\mathcal{X}}(x_{\ell}^{\mathrm{grid}})=(k(x_{1},x_{\ell}^{\mathrm{grid}}),\dots,k(x_{d},x_{\ell}^{\mathrm{grid}}))^{\top}.$
@@
$\widehat{\mathcal{K}}^{\mathrm{MC}}f(x_{\ell}^{\mathrm{grid}})=f_{\mathcal{X}}^{\top}\,\mathbf{K}_{\mathcal{X}}^{-1}\,\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}\,\mathbf{K}_{\mathcal{X}}^{-1}\,k_{\mathcal{X}}(x_{\ell}^{\mathrm{grid}}).$
@@
In the setting of the stochastic Koopman operator (3), we set $X=X_{0}$ and $Y=X_{T}$. This means that we are interested in the evolution of observables $f:\Omega\to\mathbb{R}$ under the stochastic process that solves (48), i.e.
@@
$\widehat{\mathcal{K}}^{\mathrm{MC}}f=f_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}}.$
@@
$\widehat{\mathcal{K}}^{\mathrm{MC}}f=f_{\mathcal{X}}^{\top}(\mathbf{K}_{\mathcal{X}}^{-1}+\lambda I)^{-1}\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}(\mathbf{K}_{\mathcal{X}}+\lambda I)^{-1}\mathbf{k}_{\mathcal{X}}.$
@@
    Conditional expectations of the form $\mathcal{K}f(x_{0})=\mathbb{E}[f(X_{T})\mid X_{0}=x_{0}]$ for some $x_{0}\in\Omega$ usually do not necessarily have an analytical expression. Therefore, we do not have access to the exact action of the Koopman operator on the observable $f$ (”ground truth”). Instead, we use a further Monte Carlo approximation
@@
    The randomness of the propagated matrix $K_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}$ necessitates a comparison over multiple realizations of the Koopman approximation. In our SDE setting, this means that we build matrices $\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC,n}}$, $n=1,\dots,n_{\mathrm{pred}}$ and use the mean values and standard deviations of the empirical $L^{\infty}$-errors $\widehat{\mathrm{err}}^{(n)}$ per realization to judge the performance of the approximation $\widehat{\mathcal{K}}^{\mathrm{MC}}$.
Theorem 2.6 decomposes the $L^{\infty}$-error of the stochastic Koopman approximation into a deterministic part controlled by the fill distance $h_{\mathcal{X}}$ and a probabilistic part influenced by the Monte Carlo sample size, which we denote by $m_{\mathrm{train}}$ to differentiate it from the number of Monte Carlo samples $m_{\mathrm{truth}}$ used for ground truth approximation. For numerical testing, we specify $h_{\mathcal{X}}$ (or corresponding design points $\mathcal{X}$) and $m_{\mathrm{train}}$. Throughout the testing procedure, we frequently sample the SDE solution with initial values $x_{0}\in\Omega$ at time $T>0$ from the conditional distribution $\rho_{x_{0},T}$. First, we start trajectories at chosen evaluation grid points $\mathcal{X}^{\mathrm{grid}}\subset\Omega$ for ground truth computation. Later, we repeatedly simulate trajectories starting at the design points $\mathcal{X}$ to construct realizations $\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}$ in the Koopman approximation $\widehat{\mathcal{K}}^{\mathrm{MC}}$ Algorithm 2 illustrates the procedure.
@@
For numerical stability, we apply a Tikhonov regularization with parameter $\lambda$ as in (44) to alleviate the bad conditioning of kernel matrices $\mathbf{K}_{\mathcal{X}}$, . The parameters $m_{\mathrm{train}}$ and $\lambda$ are occasionally adapted to the chosen fill distance $h_{\mathcal{X}}$ as indicated in the table of Figure 1.
@@
$\text{Wendland kernel}\quad k_{1,1}^{\text{wendland}}(x,y)$ $=(1-r)_{+}^{3}\cdot(3r+1),~\text{and}$
@@ 5.1. Linear SDE: Ornstein-Uhlenbeck process
Figure 2. Ornstein-Uhlenbeck process: mean empirical $L^{\infty}$-error with standard deviation of $n_{\mathrm{pred}}=30$ realizations of the stochastic Koopman approximation in dependence of the number of Monte Carlo simulations $m_{\mathrm{train}}$. $\star$ and $\diamond$ indicate Wendland and Matérn kernels, respectively.
@@ 5.2. Nonlinear SDE: Rescaled double-well potential
For the nonlinear SDE with double-well potential, we also set $\sigma=0.2$ and $T=1$. For better comparability, especially in terms of the fill distance, we want to restrict the corresponding SDE to the same domain $\Omega=(0,1)$. Since the wells in the regular double-well SDE (51) are at $\pm 1$, we apply the affine transformation $Z_{t}=(X_{t}+2)/4$ to mimic the behavior of the original double-well potential on (2,2)(-2,2) on our domain $\Omega=(0,1)$. Solving for $X_{t}$ gives $X_{t}=4Z_{t}-2$, so that the chain rule yields the SDE
$dZ_{t}$ $=\frac{1}{4}\left[(4Z_{t}-2)-(4Z_{t}-2)^{3}\right]\,dt\,+\frac{\sigma}{4}\,dW_{t}+dL_{t}-dU_{t}$ (55)
$Z_{0}$ $=z_{0}\in\Omega$
@@ 6. Conclusions and Outlook
We proposed a Monte Carlo-based estimator in a kernel EDMD approach for the Koopman operator of stochastic systems and analyzed its pointwise error. Extending the results of [KohnPhil24], we decompose the error into the well-known deterministic projection error and the probabilistic error introduced by the entry-wise estimation of expected values in the exact “propagation matrix” $\mathbf{K}_{\mathcal{X},\mathcal{Y}}$. To apply known projection error bounds, we derive sufficient conditions for the invariance of (fractional) Sobolev spaces $H^{s}$ under the stochastic Koopman operator, which are induced as RKHS by Wendland and Matérn kernels. Novel probabilistic error bounds for our estimator are proven to obtain a pointwise bound on the full approximation error with high probability.
@@
$\leq 2\sum_{j=1}^{d}\exp\left(-\frac{m_{j}\varepsilon^{2}}{2d\|k\|_{\infty}}\right).$
@@
$\mathbb{P}\big(\|g-g^{\mathrm{MC}}\|_{2}\geq\varepsilon\|f\|_{H})\,\leq\,2d\exp\left(-\frac{m\varepsilon^{2}}{2d\|k\|_{\infty}}\right).$
We may improve this by using the Hilbert-space Hoeffding inequality. For this, let $m_{j}=m$ for all $j$, and set $\xi_{l}=[(S_{\mathcal{X}}f)(y_{l,j})]_{j=1}^{d}-g\in\mathbb{R}^{d}$, $l\in\{1,\dots,m\}$. Then $\mathbb{E}[\xi_{l}]=0$ and
@@
Since $|(S_{\mathcal{X}}f)(y)|\leq\|k\|_{\infty}^{1/2}\|f\|_{H}$, we obtain $\|\xi_{l}\|_{2}^{2}\leq 4d\|k\|_{\infty}\|f\|_{H}^{2}$ for all $l\in\{1,\dots,m\}$. Furthermore,
$\mathbb{P}\big(\|g-g^{\mathrm{MC}}\|_{2}\geq\varepsilon\|f\|_{H})$ $=\mathbb{P}\left(\sum_{j=1}^{d}\left|\frac{1}{m}\sum_{l=1}^{m}\xi_{l,j}\right|^{2}\,\geq\,\varepsilon^{2}\|f\|_{H}^{2}\right)$
$=\mathbb{P}\left(\left\|\frac{1}{m}\sum_{l=1}^{m}\xi_{l}\right\|\,\geq\,\varepsilon\|f\|_{H}\right)$
$\leq 2\exp\left(-\frac{m\varepsilon^{2}}{8d\|k\|_{\infty}}\right).$
@@
$=\sup_{f\in V_{\mathcal{X}},\,f\neq 0}\frac{|\langle f,\Phi_{x}\rangle_{H}|}{\|f_{\mathcal{X}}\|_{2}}\,\leq\,\|k\|_{\infty}^{1/2}\sup_{f\in V_{\mathcal{X}},\,f\neq 0}\frac{\|f\|_{H}}{\|f_{\mathcal{X}}\|_{2}}=\|k\|_{\infty}^{1/2}\|\mathbf{K}_{\mathcal{X}}^{-1}\|_{2\to 2}^{1/2}.$
@@ Appendix B Recap: Wendland kernels
Wendland kernels are constructed using particular compactly supported radial basis functions, which are constructed in the following way: For some integer $m\in\mathbb{N}$, define the a function $\beta_{m}:\mathbb{R}_{+}\to\mathbb{R},\beta_{m}(r)=(1-r)_{+}^{m}=\max(1-r,0)^{m}$. Given the operator $\mathcal{I}$ that acts on measurable non-negative functions $\beta:\mathbb{R}_{+}\to\mathbb{R}_{+}$ by
$\mathcal{I}\beta(r)=\int_{r}^{\infty}t\beta(t)\,dt$
and dimension $n\geq 1$, we choose a smoothness $l\in\mathbb{N}$ and set $\beta_{n,l}=\mathcal{I}^{l}\beta_{\lfloor\frac{n}{2}\rfloor+l+1}$. This function is of the form
@@
with a univariate polynomial $p_{n,l}$ of degree $\lfloor\frac{n}{2}\rfloor+3l+1$ whose coefficients can be computed recursively by [Wend04, Theorem 9.12]. We can now define the Wendland radial basis function $B_{n,m}:\mathbb{R}^{n}\to\mathbb{R}$ of dimension $n$ and smoothness degree $m$ by
@@
and the corresponding Wendland kernel by $k_{n,l}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R},k_{n,l}(x,y)=B_{n,l}(x-y)=\beta_{n,l}(\|x-y\|_{2})$.

### v2

@@ 5. Numerical examples
$X_{0}$ $=x_{0}\in\Omega,$
with Lipschitz continuous drift $b:\mathbb{R}\to\mathbb{R}$ and diffusion $\sigma:\mathbb{R}\to\mathbb{R}$, and Brownian motion $(W_{t})_{t\geq 0}$, via the Euler-Maruyama method (see, e.g., [21]). [37, Theorem 5.2.1] guarantees the existence of a unique solution $(X_{t})_{t\geq 0}$. A basic example for this case is the Ornstein-Uhlenbeck process (OU process) given by the linear SDE
@@
We fix a timestep $T>0$, set $X=X_{0}$ and $Y=X_{T}$. To apply the setting of Sections 3 and 4, the solution process should remain in some bounded domain $\Omega=(a,b)$ until time $T$. [41] elaborates on a sufficient condition on the drift and diffusion for the invariance of a compact interval under the SDE flow. Another way to guarantee the boundedness is by reflection at the boundaries. For this, we consider the doubly reflected OU process on $\Omega=(0,1)$. It is formally given by the SDE
@@
where $L_{t}\geq 0$ and $U_{t}\geq 0$ are the minimal nondecreasing processes that ensure $X_{t}\geq 0$ and $X_{t}\leq 1$ with reflection at the boundaries for $t\geq 0$, respectively. The processes $L$ and $U$ can be interpreted as “idleness process” and “loss process”; they satisfy $\int_{0}^{T}\chi_{\{X_{t}>0\}}\,dL_{t}=0$ and $\int_{0}^{T}\chi_{\{X_{t}<1\}}\,dU_{t}=0$ for any $T>0$, respectively. We refer to [18] for details.
@@
$\mathbf{K}_{\mathcal{X}}=\bigl(\text{\sf{k}}(x_{i},x_{j})\bigr)_{i,j=1}^{d}.$
@@
$(\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}})_{ij}=\frac{1}{m_{\mathrm{train}}}\sum_{m=1}^{m_{\mathrm{train}}}\text{\sf{k}}(x_{i},y_{j}^{(m)}).$
@@
$\mathbf{k}_{\mathcal{X}}(x_{\ell}^{\mathrm{grid}})=(\text{\sf{k}}(x_{1},x_{\ell}^{\mathrm{grid}}),\dots,\text{\sf{k}}(x_{d},x_{\ell}^{\mathrm{grid}}))^{\top}.$
@@
$\widehat{\mathcal{K}}^{\mathrm{MC}}f(x_{\ell}^{\mathrm{grid}})=f_{\mathcal{X}}^{\top}\,\mathbf{K}_{\mathcal{X}}^{-1}\,\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}\,\mathbf{K}_{\mathcal{X}}^{-1}\,\mathbf{k}_{\mathcal{X}}(x_{\ell}^{\mathrm{grid}}).$
@@
In the setting of the stochastic Koopman operator (3), we set $X=X_{0}$ and $Y=X_{T}$. This means that we are interested in the evolution of observables $f:\Omega\to\mathbb{R}$ under the stochastic process that solves (49), i.e.,
@@
$\widehat{\mathcal{K}}^{\mathrm{MC}}f=f_{\mathcal{X}}^{\top}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}\mathbf{K}_{\mathcal{X}}^{-1}\mathbf{k}_{\mathcal{X}},$
@@
$\widehat{\mathcal{K}}^{\mathrm{MC}}f=f_{\mathcal{X}}^{\top}(\mathbf{K}_{\mathcal{X}}+\lambda I)^{-1}\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}(\mathbf{K}_{\mathcal{X}}+\lambda I)^{-1}\mathbf{k}_{\mathcal{X}}.$
@@
    Conditional expectations of the form $\mathcal{K}f(x_{0})=\mathbb{E}[f(X_{T})\mid X_{0}=x_{0}]$ for some $x_{0}\in\Omega$ do not necessarily admit an analytical expression. Therefore, we usually do not have access to the exact action of the Koopman operator on the observable $f$ (”ground truth”). Instead, we use a further Monte Carlo approximation
@@
    The randomness of the propagated matrix $K_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}$ necessitates a comparison over multiple realizations of the Koopman approximation. In our SDE setting, this means that we build matrices $\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC,n}}$, $n=1,\dots,n_{\mathrm{pred}}$, and use the mean values and standard deviations of the empirical $L^{\infty}$-errors $\widehat{\mathrm{err}}^{(n)}$ per realization to judge the performance of the approximation $\widehat{\mathcal{K}}^{\mathrm{MC}}$.
Theorem 2.6 decomposes the $L^{\infty}$-error of the stochastic Koopman approximation into a deterministic part controlled by the fill distance $h_{\mathcal{X}}$ and a probabilistic part influenced by the Monte Carlo sample size, which we denote by $m_{\mathrm{train}}$ to differentiate it from the number of Monte Carlo samples $m_{\mathrm{truth}}$ used for ground truth approximation. Throughout the testing procedure, we frequently sample the SDE solution with initial values $x_{0}\in\Omega$ at time $T>0$ from the conditional distribution $\rho_{x_{0},T}$. First, we start trajectories at chosen evaluation grid points $\mathcal{X}^{\mathrm{grid}}\subset\Omega$ for ground truth computation. Later, we repeatedly simulate trajectories starting at the design points $\mathcal{X}$ to construct realizations $\mathbf{K}_{\mathcal{X},\mathcal{Y}}^{\mathrm{MC}}$ in the Koopman approximation $\widehat{\mathcal{K}}^{\mathrm{MC}}$. For numerical demonstration on the domain $\Omega=(0,1)$, we first chose fill distances $h_{\mathcal{X}}$ such that $d=2/h_{\mathcal{X}}\in\mathbb{N}$. The fill distance was then achieved by spacing out the $d$ design points $\mathcal{X}=\{x_{1},\dots,x_{d}\}$ evenly on $\Omega=(0,1)$: We set $x_{1}=h_{\mathcal{X}}$, $x_{d}=1-h_{\mathcal{X}}$, and $x_{i}=(2i-1)h_{\mathcal{X}}$ for $i=2,\dots,d-1$. Algorithm 2 illustrates the procedure.
@@
For numerical stability, we apply a Tikhonov regularization with parameter $\lambda$ as in (45) to alleviate the bad conditioning of kernel matrices $\mathbf{K}_{\mathcal{X}}$. The parameters $m_{\mathrm{train}}$ and $\lambda$ are occasionally adapted to the chosen fill distance $h_{\mathcal{X}}$ as indicated in the table of Figure 1.
@@
$\text{Wendland kernel}\quad\text{\sf{k}}_{1,1}^{\text{wendland}}(x,y)$ $=(1-r)_{+}^{3}\cdot(3r+1),~\text{and}$
@@ 5.1. Linear SDE: Ornstein-Uhlenbeck process
Figure 2. Ornstein-Uhlenbeck process: mean empirical $L^{\infty}$-error of $n_{\mathrm{pred}}=30$ realizations of the stochastic Koopman approximation in dependence of the number of Monte Carlo simulations $m_{\mathrm{train}}$. $\star$ and $\diamond$ indicate Wendland and Matérn kernels, respectively.
@@ 5.2. Nonlinear SDE: Rescaled double-well potential
For the nonlinear SDE with double-well potential, we also set $\sigma=0.2$ and $T=1$. For better comparability, especially in terms of the fill distance, we want to restrict the corresponding SDE to the same domain $\Omega=(0,1)$. Since the wells in the regular double-well SDE (52) are at $\pm 1$, we apply the affine transformation $Z_{t}=(X_{t}+2)/4$ to mimic the behavior of the original double-well potential on (2,2)(-2,2) on $\Omega=(0,1)$. Solving for $X_{t}$ gives $X_{t}=4Z_{t}-2$, so that the chain rule yields the SDE
$dZ_{t}$ $=\frac{1}{4}\left[(4Z_{t}-2)-(4Z_{t}-2)^{3}\right]\,dt\,+\frac{\sigma}{4}\,dW_{t}+dL_{t}-dU_{t},$ (56)
$Z_{0}$ $=z_{0}\in\Omega,$
@@ 6. Conclusions and Outlook
We proposed a Monte Carlo-based estimator in a kernel EDMD approach for the Koopman operator of stochastic systems and analyzed its pointwise error. Extending the results of [24], we decompose the error into the well-known deterministic projection error and the probabilistic error introduced by the entry-wise estimation of expected values in the exact “propagation matrix” $\mathbf{K}_{\mathcal{X},\mathcal{Y}}$. To apply known projection error bounds, we derive sufficient conditions for the invariance of (fractional) Sobolev spaces $H^{s}(\Omega)$ under the stochastic Koopman operator, which are induced as RKHS by Wendland and Matérn kernels. Novel probabilistic error bounds for our estimator are proven to obtain a pointwise bound on the full approximation error with high probability.
@@
$\leq 2\sum_{j=1}^{d}\exp\left(-\frac{m_{j}\varepsilon^{2}}{2d\|\text{\sf{k}}\|_{\infty}}\right).$
@@
$\mathbb{P}\big(\|g-g^{\mathrm{MC}}\|_{2}\geq\varepsilon\|f\|_{H})\,\leq\,2d\exp\left(-\frac{m\varepsilon^{2}}{2d\|\text{\sf{k}}\|_{\infty}}\right).$
We may improve this by using the Hilbert-space Hoeffding inequality. For this, let $m_{j}=m$ for all $j$, and set $\xi_{l}=[(S_{\mathcal{X}}f)(y_{l,j})]_{j=1}^{d}-g\in\mathbb{R}^{d}$, $l\in\{1,\dots,m\}$. Then, we have $\mathbb{E}[\xi_{l}]=0$ and
@@
Since $|(S_{\mathcal{X}}f)(y)|\leq\|\text{\sf{k}}\|_{\infty}^{1/2}\|f\|_{H}$, we obtain $\|\xi_{l}\|_{2}^{2}\leq 4d\|\text{\sf{k}}\|_{\infty}\|f\|_{H}^{2}$ for all $l\in\{1,\dots,m\}$. Also,
$\mathbb{P}\big(\|g-g^{\mathrm{MC}}\|_{2}\geq\varepsilon\|f\|_{H})$ $=\mathbb{P}\left(\sum_{j=1}^{d}\left|\frac{1}{m}\sum_{l=1}^{m}\xi_{l,j}\right|^{2}\,\geq\,\varepsilon^{2}\|f\|_{H}^{2}\right)=\mathbb{P}\left(\left\|\frac{1}{m}\sum_{l=1}^{m}\xi_{l}\right\|\,\geq\,\varepsilon\|f\|_{H}\right)$
$\leq 2\exp\left(-\frac{m\varepsilon^{2}}{8d\|\text{\sf{k}}\|_{\infty}}\right).$
@@
$=\sup_{f\in V_{\mathcal{X}},\,f\neq 0}\frac{|\langle f,\Phi_{x}\rangle_{H}|}{\|f_{\mathcal{X}}\|_{2}}\,\leq\,\|\text{\sf{k}}\|_{\infty}^{1/2}\sup_{f\in V_{\mathcal{X}},\,f\neq 0}\frac{\|f\|_{H}}{\|f_{\mathcal{X}}\|_{2}}=\|\text{\sf{k}}\|_{\infty}^{1/2}\|\mathbf{K}_{\mathcal{X}}^{-1}\|_{2\to 2}^{1/2}.$
@@ Appendix B Recap: Wendland kernels
Wendland kernels are constructed using particular compactly supported radial basis functions, which are constructed in the following way: For $m\in\mathbb{N}$, define the function $\beta_{m}:\mathbb{R}_{+}\to\mathbb{R},\beta_{m}(r)=(1-r)_{+}^{m}{=\max(1-r,0)^{m}}$. Given the operator $\mathcal{I}$ that acts on measurable non-negative functions $\beta:\mathbb{R}_{+}\to\mathbb{R}_{+}$ by
$\mathcal{I}\beta(r)=\int_{r}^{\infty}t\beta(t)\,dt,$
and given dimension $n\geq 1$, we choose a smoothness $l\in\mathbb{N}$ and set $\beta_{n,l}=\mathcal{I}^{l}\beta_{\lfloor\frac{n}{2}\rfloor+l+1}$. This function is of the form
@@
with a univariate polynomial $p_{n,l}$ of degree $\lfloor\frac{n}{2}\rfloor+3l+1$ whose coefficients can be computed recursively by [50, Theorem 9.12]. We can now define the Wendland radial basis function $B_{n,m}:\mathbb{R}^{n}\to\mathbb{R}$ of dimension $n$ and smoothness degree $m$ by
@@
and the corresponding Wendland kernel by $\text{\sf{k}}_{n,l}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R},\text{\sf{k}}_{n,l}(x,y)=B_{n,l}(x-y)=\beta_{n,l}(\|x-y\|_{2})$.